f(x)=x^5+2x^3+2x  how to find out inverse of f(x)?

Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear barnita,

Finding the inverse of a function is quite easy.

f(x)=x^5+2x^3+2x

y=x^5+2x^3+2x

y/x=x^4+2x^2+2

x=y/(x^4+2x^2+2)

Hence the inverse is: x/(y^4+2y^2+2)

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

 

Sagar Singh

B.Tech IIT Delhi